# AIM: Adaptive Inference of Multi-Modal LLMs via Token Merging and Pruning

### v1

@@ Abstract
Large language models (LLMs) have enabled the creation of multi-modal LLMs that exhibit strong comprehension of visual data such as images and videos. However, these models usually rely on extensive visual tokens from visual encoders, leading to high computational demands, which limits their applicability in resource-constrained environments and for long-context tasks. In this work, we propose a training-free adaptive inference method for multi-modal LLMs that can accommodate a broad range of efficiency requirements with a minimum performance drop. Our method consists of a) iterative token merging based on embedding similarity before LLMs, and b) progressive token pruning within LLM layers based on multi-modal importance. With a minimalist design, our method can be applied to both video and image LLMs. Extensive experiments on diverse video and image benchmarks demonstrate that, our method substantially reduces computation load (e.g., a 7-fold reduction in FLOPs) while preserving the performance of video and image LLMs. Further, under a similar computational cost, our method outperforms the state-of-the-art methods in long video understanding (e.g., +4.6 on MLVU). Additionally, our in-depth analysis provides insights into token redundancy and LLM layer behaviors, offering guidance for future research in designing efficient multi-modal LLMs. Our code will be available at https://github.com/LaVi-Lab/AIM.
@@ 1 Introduction
Large language models (LLMs) [52, 89, 63, 6, 80] have been recently adapted for visual understanding, fostering the developments in both image [38, 7, 96, 4, 22] and video LLMs [44, 36, 33, 77, 91]. However, these multi-modal LLMs usually rely on a large number of visual tokens generated by visual encoders [56, 86], especially for video data where token counts can reach thousands per video. Such high token number requires extensive computational resources for both training and inference, restricting their use in real-world applications (e.g., real-time processing on mobile devices). Further, as the number of video frames increases, the total number of tokens also grows, limiting the model’s capacity to handle dense video frames. This often results in the loss of crucial temporal information, which is essential for comprehending long videos [94, 88, 60].
Figure 1: Our proposed training-free method enables adaptive inference of pre-trained multi-modal LLMs, supporting a wide range of computational conditions. Compared with the base pre-trained model, our method substantially reduces FLOPs, without or with a manageable performance drop.
To bridge the gap, we propose leveraging the inherent redundancy present in visual data to develop adaptive inference for multi-modal LLMs. Adaptive inference dynamically adjusts a model’s computational load during inference based on contextual factors [19], e.g., image content, computational constraints, or desired accuracy levels. Since retaining all visual tokens during inference is often unnecessary, our key insight is to strategically select these tokens throughout the inference process. This approach allows for controlling the amount of computation necessary for inference, so as to optimize the model’s efficiency while preserving its accuracy.
In this work, we propose a training-free method of adaptive inference tailored for multi-modal LLMs, consisting of token merging based on visual similarity and token pruning based on multi-modal importance. Specifically, the visual encoder transforms the input visual data into visual tokens, which are iteratively merged based on their embedding similarities. The merged tokens are then fed into the LLM, where the tokens considered less important for multi-modal reasoning are progressively pruned at each layer, resulting in a gradual reduction of tokens. By adjusting the key parameters in the token merging and token pruning process, our method enables adaptive inference that achieves various levels of computation reduction with manageable performance loss, shown in Figure 1. With this capability, we can select a configuration of token merging and pruning so that the model meets the resource constraint while the performance is optimized.
@@ 2 Related Work
Multi-modal LLMs. Large language models [50, 51, 46] trained on large-scale datasets exhibit strong performance in text understanding and generation tasks, laying the foundation for developing multi-modal LLMs with visual instruction tuning [39, 7, 96]. To further enhance model capabilities, recent work has focused on increasing visual input resolution  [38, 40] and creating higher-quality instruction-tuning datasets  [30]. Meanwhile, advancements in video-based LLMs have also been achieved by extending visual instruction tuning to video modality. For example, VideoChatGPT [44] pioneers video-based instruction tuning through the creation of video question-answer pairs, while Video-LLaVA [36] improves model performance by employing text-aligned video features [95]. In parallel, several models have extended their capabilities from static images to dynamic video inputs, such as LLaVA-OneVision [30, 92], Qwen2-VL [66], and mPlug-Owl3 [82].
Recent work has focused on improving video LLMs for long video understanding. Given the challenge posed by the large number of tokens in long videos, recent approaches [70, 10, 49, 73, 68, 29, 88] adopt Q-former [32] or state space models [18] to aggregate the information before feeding visual tokens into the LLMs. Others extend the context window of LLMs [71, 88, 21] or design sequence parallelism from a systems perspective [79] to process the long video sequence. Built on top of the pre-trained image and video LLMs, our adaptive inference method aims to reduce token redundancy during inference and adapt to diverse efficiency requirements while maintaining model performance.
Figure 2: An overview of our training-free method for pre-trained multi-modal LLMs, optimized for accuracy-efficiency trade-off. During inference, we first merge the input visual tokens of LLM based on the cosine similarity between token embeddings, reducing their redundancy. The retained tokens are then fed into the LLM. Token pruning is applied within the LLM layers using the Page Rank algorithm, with a scheduler controlling the retention ratio of each layer. By adjusting the merging ratio and pruning scheduler parameters, our approach enables adaptive inference of multi-modal LLMs, accommodating various computational constraints with minimal performance loss.
@@
For multi-modal LLMs, token pruning has been explored recently [5, 59]. FastV [5] prunes visual tokens at a particular selected LLM layer, while LLaVA-Prumerge [59] leverages the key-value pairs from the vision encoder to prune visual tokens before LLM. Unlike them, our method performs token reduction both before and across LLM layers, allowing for adaptive inference fitting with various computation constraints.
There are concurrent efforts on arXiv [93, 76, 83, 90] that also seek to reduce the computation of multi-modal LLMs. They are only applied to image LLMs, prune tokens exclusively within LLM, or rely on fine-tuning for an accuracy-efficiency balance. In comparison, our work is training-free, prunes visual tokens before and within LLM, generalizes to both video and image LLMs, and moreover, supports adaptive inference that accommodates a wide range of computation budgets with minimal performance loss.
Adaptive Inference. The ability to dynamically adjust the computational complexity of prediction models based on input data, latency constraints, or desired accuracy levels has received considerable interest in the vision and learning community [19]. Early attempts primarily focus on feature selection within multi-stage prediction pipelines [25, 78, 17]. More recent efforts have extended this concept to the adaptive inference of deep models. Methods have been developed for both convolutional networks and vision Transformers, enabling techniques such as input resampling, operation skipping, layer dropping, and early exiting [12, 31, 67, 2, 74, 20, 24, 48, 69, 57, 54, 47]. Similar ideas have also been explored for LLMs [9, 58]. In contrast to previous work, we propose an adaptive inference method tailored for multi-modal LLMs via merging similar tokens and pruning unimportant tokens in multi-modal reasoning.
@@ 3.1 Multi-modal LLMs
Our training-free method is designed to reduce the redundancy of visual tokens, by applying it to the input visual tokens of the LLM and to the visual tokens at each LLM layer. We denote the visual tokens right before LLM as $\mathbf{v^{0}}\in\mathcal{R}^{N_{0}\times D}$ where $N_{0}$ is the initial number of visual tokens. Within LLM, we denote the input visual tokens and the input text tokens at the $l$-th layer as $\mathbf{v^{l}}\in\mathcal{R}^{N_{l}\times D}$ and $\mathbf{t^{l}}\in\mathcal{R}^{M_{l}\times D}$, respectively. $l\in[1,L]$ denotes the index of LLM layer and $N_{l}$ ($M_{l}$) represents the number of visual (text) tokens at layer $l$.
@@ 3.2 Token Merging before LLM
The input to LLMs often consists of hundreds of visual tokens for an image or thousands for a video, resulting in substantial redundancy. Merging these redundant tokens before they enter the LLM can significantly reduce computational demands. Inspired by [3], we merge the visual tokens with high similarity, where similarity is quantified by the cosine similarity between token embeddings. As illustrated in Figure 2, given an initial set of visual tokens $\mathbf{v^{0}}$ at the first LLM layer, we divide adjacent tokens into sets A and B, calculate pairwise similarity scores between the sets, and identify the closest matching token in set B for each token in set A. We then merge the token pairs with the highest similarity scores by averaging their embeddings. This process reduces the number of tokens by half at most. We repeat this merging process iteratively (e.g., twice) to achieve the desired retention ratio.
For videos, we merge the visual tokens within individual frames. Empirical studies suggest that merging tokens within individual frames has minimal impact on final reasoning performance in video tasks, while merging across frames can be harmful. We hypothesize that merging across frames disrupts temporal order of tokens, leading to a loss of crucial temporal information for video understanding.
@@ 3.3 Token Pruning within LLM
where $s^{l}_{j}$ is initialized uniformly over all tokens and $\mathbf{A}^{l}$ represents the softmax-normalized attention weights. Based on these importance scores, we only prune visual tokens and retain the most important visual tokens, leaving text tokens intact. Our rationale is that pruning text tokens substantially degrades performance, likely because multi-modal LLMs rely on text tokens to perform text-centric reasoning. As a result, the number of visual tokens input to the next layer is $N_{0}\times r^{l}$, where $r^{l}$ is the retention ratio at layer $l$.
@@
$r_{l}=\small\begin{cases}1,&\text{if }l<l_{1}\\ 1-k(l-l_{1}),&\text{if }l_{1}\leq l\leq l_{2}\\ 0,&\text{if }l>l_{2}\end{cases}$ (2)
@@ 3.4 Adaptive Inference
Table 1: Performance on video benchmarks. Compared to the base model LLaVA-OV-7B, our method significantly reduces FLOPs and prefill time by a factor of 6.8 and 8.0, respectively, with minimum or even without performance loss. Compared to baseline methods, our method outperforms them on most benchmarks with only 62.4% FLOPs and 63.3% prefill time.
4 Experiments
In this section, we first introduce our implementation details and benchmarks, and then present our results, including the results of video benchmarks, image benchmarks, ablation studies, and adaptive inference.
@@ 4 Experiments
Table 2: Performance on long video benchmarks with more sampled frames as inputs. With comparable FLOPs and prefill time, our method can accommodate more sampled frames and thus improve the base model LLaVA-OV-7B for long video understanding.
@@ 4.1 Video Benchmarks
Baselines. We choose baselines as FastV [5] which prunes visual tokens at a particular LLM layer, and LLaVA-1.5-Prumerge [59] which leverages the key-query pairs in the visual encoder to prune and merge visual tokens. We run their official implementation with the pre-trained LLaVA-OV-7B model in the training-free setting.
Accuracy-efficiency balance. As Table 1 shows, our model achieves substantial reductions in computational demands compared to the base model LLaVA-OV-7b (e.g., 14.76 vs. 99.63 FLOPs), decreasing FLOPs by a factor of 6.8 and prefill time by 8.0, with little or no performance drop across diverse benchmarks. Additionally, compared to the baseline models FastV and LLaVA-Prumerge, our model consistently achieves higher performance on most benchmarks, while only requiring 62.4% FLOPs and 63.3% prefill time at most. These results indicate that our method effectively reduces redundancy in visual tokens, providing an optimal balance between accuracy and efficiency.
@@
Table 3: Performance on image benchmarks. VQA-v2 and GQA are the most stable benchmarks that have the most evaluation samples. Compared to the base model LLaVA-1.5-7B, our model significantly reduces FLOPs and prefill time (e.g., by a factor of 3.7 and 2.7, respectively), with a small performance loss. With the same level of FLOPs and prefill time, our method outperforms the baseline methods by a clear margin on most benchmarks.
@@ 4.2 Image Benchmarks
The results are reported in Table 3, comparing our model with the base model and baseline method.
@@
Baselines. We again choose the baseline as FastV [5] and LLaVA-1.5-Prumerge [59] which provide training-free results on image benchmarks. We also compare with a variant LLaVA-1.5-Prumerge+ which trades computation for reasoning performance. Both our models and baseline models are evaluated in the training-free setting with the pre-trained LLaVA-1.5-7B as the base model.
@@
Comparison with baseline methods. In Table 3, with the same level of FLOPs and prefill time, our model consistently outperforms the baseline models on most benchmarks by a clear margin (e.g., +2.3 on VQA-v2, +3.3 on GQA, +33.2 on MME, +2.3 on MMB, +3.3 on POPE over LLaVA-Prumerge). These results suggest that our method can effectively retain visual tokens critical for image reasoning tasks, while reducing computation costs.
@@ 4.3 Ablation Study
In this section, we conduct ablation studies for our method, by adjusting the retention ratio of token merging, the parameters of the token pruning scheduler, and the pruning strategy. Their results are presented in Table 4, Table 5, and Table 6, respectively.
@@
Table 4: Ablation study on token merging. We show the performance on the VideoMME benchmark by varying various retention ratios of token merging, with token pruning within LLMs disabled.
Accuracy-efficiency trade-offs controlled by token pruning scheduler. As shown in Table 5, by adjusting the pruning scheduler parameters $l_{1}$ and $l_{2}$, our method enables a broad range of accuracy-efficiency trade-offs. FLOPs decrease from 22.90 to 6.71, and prefill time drops from 83.94 to 28.18, with performance unaffected until FLOPs and prefill time are reduced to approximately half. Furthermore, in each block, when comparing the first row with subsequent rows, we observe a reduction in FLOPs and prefill time with little to no impact on performance. For example, Exp. 6 achieves a 16% reduction in both FLOPs and prefill time compared to Exp. 5, while maintaining performance. These findings validate our token pruning design, demonstrating its flexibility in controlling LLM computations and achieving optimal accuracy-efficiency trade-offs.
Early LLM layers vs. later LLM layers. We find that LLM emphasizes multi-modal fusion at early layers and shifts the focus to text-centric reasoning at later layers. For example, experiments 5, 8, and 10 in Table 5 remove all visual tokens after layer 22, layer 15, and layer 8, respectively. Performance is maintained when visual tokens are removed starting from layer 22 (Exp. 5), but it declines when tokens are removed from layer 15 (Exp. 8) and more significantly from layer 8 (Exp. 10). Notably, removing tokens as early as layer 8 causes a substantial performance drop (e.g., from 58.0 to 41.9). Based on these findings, we adopt a strategy (Exp. 16) that retains visual tokens in early layers, gradually prunes them in middle layers, and completely removes them in later layers.
@@
Table 5: Ablation study on token pruning. We show the performance on the VideoMME benchmark by varying the parameters of our pruning scheduler ($l_{1}$ $l_{2}$), with token merging set to 25% retention ratio.
@@
Table 6: Ablation study on pruning text tokens within LLM. We show the performance on the VideoMME benchmark, with the only difference as whether our method prunes text tokens or not.

### v2

@@ Abstract
Large language models (LLMs) have enabled the creation of multi-modal LLMs that exhibit strong comprehension of visual data such as images and videos. However, these models usually rely on extensive visual tokens from visual encoders, leading to high computational demands, which limits their applicability in resource-constrained environments and for long-context tasks. In this work, we propose a training-free adaptive inference method for multi-modal LLMs that can accommodate a broad range of efficiency requirements with a minimum performance drop. Our method consists of a) iterative token merging based on embedding similarity before LLMs, and b) progressive token pruning within LLM layers based on multi-modal importance. With a minimalist design, our method can be applied to both video and image LLMs. Extensive experiments on diverse video and image benchmarks demonstrate that our method substantially reduces computation load (e.g., a 7-fold reduction in FLOPs) while preserving the performance of video and image LLMs. Further, at a similar computational cost, our method outperforms the state-of-the-art methods in long video understanding (e.g., +4.6 on MLVU). Additionally, our in-depth analysis provides insights into token redundancy and LLM layer behaviors, offering guidance for future research in designing efficient multi-modal LLMs. Our code is available at https://github.com/LaVi-Lab/AIM.
@@ 1 Introduction
Large language models (LLMs) [54, 93, 66, 6, 85] have been recently adapted for visual understanding, fostering the developments in both image [41, 8, 102, 4, 25] and video LLMs [45, 39, 36, 80, 94]. However, these multi-modal LLMs (MLLMs) usually rely on a large number of visual tokens generated by visual encoders [58, 91], especially for video data where token counts can reach thousands per video. Such high token number requires extensive computational resources for both training and inference, restricting their use in real-world applications (e.g., real-time processing on mobile devices). Further, as the number of video frames increases, the total number of tokens also grows, limiting the model’s capacity to handle dense video frames. This often results in the loss of crucial temporal information, which is essential for comprehending long videos [100, 92, 61].
Figure 1: Key idea and finding. Our training-free method enables adaptive inference of pre-trained multi-modal LLMs, supporting a wide range of computational conditions. In comparison to the base pre-trained model, our method substantially reduces FLOPs, without or with a manageable performance drop.
To bridge the gap, we propose leveraging the inherent redundancy present in visual data to develop adaptive inference for multi-modal LLMs. Adaptive inference dynamically adjusts a model’s computational load during inference based on contextual factors [22], e.g., computational constraints, image content, or desired accuracy levels. Since retaining all visual tokens during inference is often unnecessary, our key insight is to strategically select these tokens throughout the inference process. This approach allows for controlling the amount of computation necessary for inference, so as to optimize the model’s efficiency while preserving its accuracy.
In this work, we introduce a training-free method of adaptive inference tailored for multi-modal LLMs, consisting of token merging based on visual similarity and token pruning based on multi-modal importance. Specifically, the visual encoder transforms the input visual data into visual tokens, which are iteratively merged based on their embedding similarities. The merged tokens are then fed into the LLM, where the tokens considered less important for multi-modal reasoning are progressively pruned at each layer, resulting in a gradual reduction of tokens. By adjusting the key parameters in the token merging and token pruning process, our method enables adaptive inference that achieves various levels of computation reduction with manageable performance loss, shown in Figure 1. With this capability, we can select a configuration of token merging and pruning so that the model meets the resource constraint while the performance is optimized.
@@ 2 Related Work
Multi-modal LLMs. Large language models [52, 53, 47] exhibit strong performance in text understanding and generation tasks, laying the foundation for developing image LLMs [8, 102, 41, 33] and their chain-of-thoughts [50, 99, 97]. Meanwhile, video LLMs have also made progress via video instruction tuning [45, 39, 101, 33, 95, 69, 87]. Recent works [73, 13, 51, 76, 71, 32, 92] focus on improving video LLMs for long video understanding, addressing the challenge of a large number of video frames. They either adopt Q-former [35] or state-space models [21] to aggregate the information before feeding visual tokens into the LLMs. Others extend the context window of LLMs [74, 92, 24] or design sequence parallelism from a systems perspective [84] to process the long video sequence. Built on top of the pre-trained image and video LLMs, our adaptive inference method aims to reduce token redundancy during inference and adapt to diverse efficiency requirements while maintaining model performance.
Figure 2: Overview of our training-free method for pre-trained multi-modal LLMs, optimized for accuracy-efficiency trade-off. During inference, we first merge the input visual tokens of LLM based on the cosine similarity between token embeddings, reducing their redundancy. The retained tokens are then fed into the LLM. Token pruning is applied within the LLM layers using the Page Rank algorithm, with a scheduler controlling the retention ratio of each layer. By adjusting the merging ratio and pruning scheduler parameters, our approach enables adaptive inference of multi-modal LLMs, accommodating various computational constraints with minimal performance loss.
@@
For multi-modal LLMs, token pruning has been explored recently [5, 40, 79, 62]. FastV [5] and VTW [40] prune visual tokens at a particular selected LLM layer, while PDrop [79] prunes tokens at the end of each stage of LLM layers. LLaVA-Prumerge [62] leverages the key-value pairs from the vision encoder to prune visual tokens before LLM. Unlike them, our method performs token reduction both before and across LLM layers, allowing for adaptive inference fitting with various computation constraints.
There are concurrent efforts [98, 88, 96, 65] that also seek to reduce the computation of multi-modal LLMs. They are only applied to image LLMs, prune tokens exclusively within LLM, or rely on fine-tuning for an accuracy-efficiency balance. In comparison, our work is training-free, prunes visual tokens before and within LLM, generalizes to both video and image LLMs, and moreover, supports adaptive inference that accommodates a wide range of computation budgets with minimal performance loss.
Adaptive Inference. The ability to dynamically adjust the computational complexity of prediction models based on input data, latency constraints, or desired accuracy levels has received considerable interest in the vision and learning community [22]. Early attempts mainly focus on feature selection within multistage prediction pipelines [28, 82, 20]. More recent efforts have extended this concept to the adaptive inference of deep models. Methods have been developed for both convolutional networks and vision Transformers, enabling techniques such as input resampling, operation skipping, layer dropping, and early exiting [15, 34, 70, 2, 77, 23, 27, 49, 72, 59, 56, 48, 81]. Similar ideas have also been explored for LLMs [12, 60], and recently extended to MLLMs [83]. In contrast to these works, we propose an adaptive inference method tailored for multi-modal LLMs via merging similar tokens and pruning unimportant tokens in multi-modal reasoning.
@@ 3.1 Multi-modal LLMs
Our training-free method is designed to reduce the redundancy of visual tokens, by applying it to the input visual tokens of the LLM and to the visual tokens at each LLM layer. We denote the visual tokens right before LLM as $\mathbf{v^{0}}\in\mathcal{R}^{N^{0}\times D}$, where $N^{0}$ is the initial number of visual tokens. Within LLM, we denote the input visual tokens and the input text tokens at the $l$-th layer as $\mathbf{v^{l}}\in\mathcal{R}^{N^{l}\times D}$ and $\mathbf{t^{l}}\in\mathcal{R}^{M^{l}\times D}$, respectively. $l\in[1,L]$ denotes the index of LLM layer and $N^{l}$ ($M^{l}$) represents the number of visual (text) tokens at layer $l$.
@@ 3.2 Token Merging before LLM
The input to LLMs often consists of hundreds of visual tokens for an image or thousands for a video, resulting in substantial redundancy. Merging these redundant tokens before they enter the LLM can significantly reduce computational demands. Inspired by [3], we merge the visual tokens with high similarity, where similarity is quantified by the cosine similarity between token embeddings. Unlike  [3] that merges tokens at each layer of visual encoder, our method perform token merging after visual encoder. This design is agnostic to encoder architectures and easy for plug-and-play. As illustrated in Figure 2, given an initial set of visual tokens $\mathbf{v^{0}}$ at the first LLM layer, we divide adjacent tokens into sets A and B, calculate pairwise similarity scores between the sets, and identify the closest matching token in set B for each token in set A. We then merge the token pairs with the highest similarity scores by averaging their embeddings. This process reduces the number of tokens by half at most. We repeat this merging process iteratively (e.g., twice) to achieve the desired retention ratio.
For videos, we merge the visual tokens within individual frames. Empirical ablation studies suggest that merging tokens within individual frames has minimal impact on final reasoning performance in video tasks, while merging across frames can be harmful. We hypothesize that merging across frames disrupts temporal order of tokens, leading to a loss of crucial temporal information for video understanding.
@@ 3.3 Token Pruning within LLM
where $s^{l}_{j}$ is initialized uniformly over all tokens and $\mathbf{A}^{l}$ represents the softmax-normalized attention weights. Based on these importance scores, we only prune visual tokens and retain the most important visual tokens, leaving text tokens intact. Our rationale is that pruning text tokens substantially degrades performance, likely because multi-modal LLMs rely on text tokens to perform text-centric reasoning. As a result, the number of visual tokens input to the next layer is $N^{1}\times r^{l}$, where $r^{l}$ is the retention ratio at layer $l$.
@@
$r^{l}=\small\begin{cases}1,&\text{if }l<l_{1}\\ 1-k(l-l_{1}),&\text{if }l_{1}\leq l\leq l_{2}\\ 0,&\text{if }l>l_{2}\end{cases}$ (2)
@@ 3.4 Adaptive Inference
Table 1: Results on video benchmarks. Compared to the base model LLaVA-OV-7B, our method significantly reduces FLOPs and prefill time by a factor of 6.8 and 8.0, respectively, with minimum or even without performance loss. Compared to baseline methods, our method outperforms them on most benchmarks with only 69.5% FLOPs and 69.2% prefill time.
@@
Table 2: Results on long video benchmarks using more sampled frames as inputs. With comparable FLOPs and prefill time, our method can accommodate more sampled frames and thus improve the base model LLaVA-OV-7B for long video understanding.
4 Experiments
In this section, we first introduce our implementation details and benchmarks, and then present our results, including the results of video benchmarks, image benchmarks, ablation studies, computation overhead, and adaptive inference.
@@ 4.1 Video Benchmarks
Table 3: Results on image benchmarks. VQA-v2 and GQA are the most stable benchmarks that have the most evaluation samples. With less computation cost, our method outperforms baselines on most benchmarks. Note that PDrop keeps all tokens at the first 25% LLM layers and thus does not support low FLOPs, i.e., below 25% FLOPs of base model. Compared to the base model LLaVA-1.5-7B, our model significantly reduces FLOPs and prefill time (e.g., by a factor of 3.7 and 2.7, respectively), with a manageable performance loss.
Baselines. Our baselines include (1) FastV [5] that prunes visual tokens at a particular LLM layer, (2) VTW [40] which abandons all visual tokens at a particular selected LLM layer, (3) PDrop [79] that divides LLM layers into four stages and prunes tokens at the end of each stage, based on the attention scores between visual tokens and the last text token, and (4) LLaVA-1.5-Prumerge [62] that leverages the key-query pairs in the visual encoder to prune and merge visual tokens. We conduct the experiments with the pre-trained LLaVA-OV-7B model in the training-free setting.
Accuracy-efficiency balance. As Table 1 shows, our model achieves substantial reductions in computational demands compared to the base model LLaVA-OV-7b (e.g., 14.76 vs. 99.63 FLOPs), decreasing FLOPs by a factor of 6.8 and prefill time by 8.0, with little or no performance drop across diverse benchmarks. Additionally, compared to the baseline methods, our model consistently achieves higher performance on most benchmarks, while only requiring 69.5% FLOPs and 69.2% prefill time at most. These results indicate that our method effectively reduces redundancy in visual tokens, providing an optimal balance between accuracy and efficiency.
@@ 4.2 Image Benchmarks
The results are reported in Table 3, comparing our model with the base model and baseline methods.
@@
Baselines. We again choose the baselines as FastV [5], VTW [40], PDrop [79], and LLaVA-1.5-Prumerge [62] which provide training-free results on image benchmarks. We also compare with a variant LLaVA-1.5-Prumerge+ which trades computation for reasoning performance. Both our models and baselines are evaluated in the training-free setting with LLaVA-1.5-7B as the base model.
@@
Comparison with baseline methods. In Table 3, with the same level of FLOPs and prefill time, our model consistently outperforms the baseline methods on most benchmarks by a clear margin (e.g., +2.3 on VQA-v2, +3.3 on GQA, +33.2 on MME, +2.3 on MMB, +3.3 on POPE over LLaVA-Prumerge). These results suggest that our method can effectively retain visual tokens critical for image reasoning tasks, while reducing computation costs.
@@ 4.3 Ablation Study
In this section, we conduct ablation studies for our method, by adjusting the retention ratio of token merging, the parameters of the token pruning scheduler, and the pruning strategy. Their results are presented in Table 4, Table 5, and Table 6, respectively. More ablations are in the Appendices.
@@
Table 4: Ablation study on token merging. We show the performance on the VideoMME benchmark by varying various retention ratios of token merging, with token pruning in LLMs disabled.
@@
Table 5: Ablation study on token pruning. We show the performance on VideoMME by varying parameters of our pruning scheduler ($l_{1}$ $l_{2}$), with 25% retention ratio for token merging.
Controlling accuracy-efficiency trade-offs with token pruning scheduler. As shown in Table 5, by adjusting the pruning scheduler parameters $l_{1}$ and $l_{2}$, our method enables a broad range of accuracy-efficiency trade-offs. FLOPs decrease from 22.90 to 6.71, and prefill time drops from 83.94 to 28.18, with performance unaffected until FLOPs and prefill time are reduced to approximately half. Furthermore, in each block, when comparing the first row with subsequent rows, we observe a reduction in FLOPs and prefill time with little to no impact on performance. For example, Exp. 6 achieves a 16% reduction in both FLOPs and prefill time compared to Exp. 5, while maintaining performance. These findings validate our token pruning design, demonstrating its flexibility in controlling LLM computations and achieving optimal accuracy-efficiency trade-offs.
Early LLM layers vs. later LLM layers. We find that LLM emphasizes multi-modal fusion at early layers and shifts the focus to text-centric reasoning at later layers. For example, experiments 5, 8, and 10 in Table 5 remove all visual tokens after layer 22, layer 15, and layer 8, respectively. Performance is maintained when visual tokens are removed starting from layer 22 (Exp. 5), but it declines when tokens are removed from layer 15 (Exp. 8) and more significantly from layer 8 (Exp. 10). Notably, removing tokens as early as layer 8 causes a substantial performance drop (e.g., from 58.0 to 41.9). Based on these findings, we adopt a strategy (Exp. 6) that retains visual tokens in early layers, gradually prunes them in middle layers, and completely removes them in later layers.
@@
Table 6: Ablation study on pruning text tokens within LLM. We report the performance on VideoMME by comparing our method with and without text token pruning.
Overhead of our method. As in Table 7, compared to the FLOPs of LLM inference, the additional FLOPs introduced by our method are minimal—only 0.6% FLOPs of Qwen2 and 0.03% FLOPs of Vicuna-v1.5. These results show that our overhead (cost) is negligible in comparison to the multi-fold FLOPs reduction (benefit).
FLOPs Video LLM Image LLM
(GB) (Qwen2-7B) (Vicuna-v1.5-7B)
Token Merging 88.25 0.23
Token Pruning 4.18 0.03
Total 92.43 0.26
LLM Inference 14757 1003
Table 7: Computational cost in GFLOPs introduced by our method. The cost is much less than the FLOPs of LLM inference.
4.4 Adaptive Inference
As shown in Table 8, by combining token merging and token pruning, our method supports a broad range of accuracy-efficiency trade-offs. For instance, compared to the base model (99.63 FLOPs and 58.2 accuracy), our default configuration achieves the same accuracy with significantly reduced FLOPs (14.76), while an efficient configuration reduces FLOPs even further (2.51) with an acceptable performance decrease (50.9 accuracy). Overall, our approach spans a 40-fold reduction in FLOPs with less than a 13% drop in accuracy. This flexibility is enabled by adjusting the key parameters, including the retention ratio in token merging and $l_{1}$ $l_{2}$ in token pruning. These configurations make our adaptive inference method suitable for various devices and efficiency requirements, such as AR glasses, mobile phones, personal computers, and robots. More results can be found in the Appendices.